# Superuniversal Statistics with Topological Origins for non-Hermitian Scattering Singularities

### v1

@@ Appendix B Singularities in $d=3$ Parameter Space
As stated in Sec. I, while singularities form speckle patterns of points in $d=2$ dimensional parameter space, in $d=3$ they form continuous curves because singularities always have dimension $d-2$. These curves either form closed loops or trail off the edge of the explored parameter space, with no end points visible. Note that this is the same property we see for the $d-1$ dimensional curves in a $d=2$ dimensional parameter space. In $d=3$ space, the $d-1$ dimensional topologically protected structures such as $\text{Re}[\text{det}S]=0$ take the form of surfaces that are either closed or extend beyond the limits of the explored parameter space. In simulations, it can be seen that sometimes $d-1$ structures in $d=2$ space extend out to infinite value of certain parameters (such as loss $\eta$), so we assume the same must be true for $d-2$ structures in higher dimensional spaces. Experimentally, our tunable perturbations are limited within some range, so we cannot verify this assumption.
@@
In Fig. 8, there is a color gradient in frequency but visually every line appears to be a single solid color. Most scattering singularities, CPAs included, do not move much in frequency under perturbations to a graph system in the form of bond length changes. In $\mathcal{D}=2$ and $3$ dimensional cavities there are more pronounced movements in frequency, though still nothing dramatic. Practically, this means it is easier to move multiple different S-matrix zeros to different spots along the real frequency axis than to drag a single zeros to different spots along the real frequency axis. For graphs, most frequency movement is on the order of the mean mode spacing or less, which for this graph is about $40$ MHz. An example of this can be seen in Fig. 4(a) of Ref. [71], which shows a CPA singularity line extending over $60$ MHz, the data coming from the same graph used in Fig. 8.
Appendix C Singularity Statistics - Random Matrix Theory
The process used to arrive at statistical results from RMT is as follows. First, we create two sets of $600$ independent Hamiltonians that are of dimension $10^{5}\crossproduct 10^{5}$, generated using GOE (for $\beta=1$) and Gaussian Unitary Ensemble (GUE) (for $\beta=2$) algorithms. Then normalized impedance matrices $Z$ are calculated using the Random Coupling Model [125, 136, 137, 53] with the only parameter changed between different ensembles being the dimensionless uniform attenuation $\eta$. The impedance matrices are converted to scattering matrices through the formula:
@@ Appendix C Singularity Statistics - Random Matrix Theory
Note that because the impedance matrices created through RCM are normalized, the characteristic impedance $Z_{0}=1\Omega$. The result is a collection of ensembles of $600$ independent $S$-matrices, each ensemble characterized by $M,\beta$ and $\eta$. We treat these scattering matrices generated through RMT the same as our experimental ensembles and conduct the same analysis to calculate singularity mode density, shown in Fig. 4(b).
Figure 9: (a-b) PDFs of $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$ and $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ from 51 RMT ensembles with different values for the three parameters $M,\beta,\eta$. Color gradient of curves corresponds to $\eta$ (not uniformaly spaced) of the ensembles. Dashed black lines characterize power law of tail behavior. (c-d) PDFs of $\frac{\langle|\text{Re}[\mathcal{S}]|\rangle}{|\text{Re}[\mathcal{S}]|}$ and $\frac{\langle|\mathcal{S}|\rangle}{|\mathcal{S}|}$ for various functions $\mathcal{S}$ from two arbitrary RMT ensembles. Dashed black lines characterize power law of tail behavior.
@@
The RMT data can also be used to calculate PDFs of various complex scalar functions. In Fig. 9, we show the same PDFs as in Fig. 3: (a) $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$, (b) $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$, and the same eight and ten functions in panels (c) and (d), respectively. The simulations utilize two sets of $600$ independent Hamiltonians that are of dimension $10^{5}\crossproduct 10^{5}$, generated using GOE (for $\beta=1$) and Gaussian Unitary Ensemble (GUE) (for $\beta=2$) algorithms. Then the scattering matrices are calculated using the Random Coupling Model [125, 136, 137, 53] with the only parameter changed between different ensembles being the dimensionless uniform attenuation $\eta$.
@@
Identical behavior is seen in Fig. 9 as in Fig. 3: $-2$ power law tails are seen panels (a) and (c), and $-3$ power law tails are seen in panels (b) and (d). This shows that even though the data in this paper is measured from microwave cavities, the results we present are true for all chaotic wave scattering systems. The present paper also explains why in the Appendix of Ref. [53], a $-2$ power law tail was seen for the probability distributions of complex reflection and transmission time delay differences ($\tau_{\delta R}$ and $\tau_{\delta T}$) from Hermitian ($\eta=0$) systems. As shown in Fig. 5(a), in a Hermitian system, $\delta R_{12}=0$ is $d-1$ dimensional, not a $d-2$ dimensional singularity. The same is true of $\delta T_{12}=0$ if the system is also non-reciprocal. Further the PDFs of complex transmission time delay from a Hermitian, reciprocal system would have $-2$ power laws, which is why the authors of Ref. [53] chose to show plots from $\beta=2$ RMT data, since at the time they did not have the topology argument to understand why they did not see $-3$ power laws.
@@
A final comment is that the PDF of $\frac{\langle|C|\rangle}{|C|}$, where $|C|$ is the coalescence of the $S$-matrix eigenvectors, from the reciprocal system shown in panel (c) has the expected $-2$ power law tail out to the largest values, never transitioning to $-3$. This is because RMT numerics lacks the small amount of residual non-reciprocity we always see in experimental data.
Appendix D Exceptional Points at Infinitesimal Absorption
@@
Figure 10: Phase of $S_{+i}$ and $S_{-i}$ from same kind of simulation as in Fig. 5 over the same parameter space, but with $\eta=6*10^{-9}$. Black circles surround EPs (phase singularities) and arrows show direction of phase winding.
@@
Appendix E Complex Time Delay - Generalizations
Every singularity can be associated with the divergence of a complex time delay. Examples are Wigner-Smith time delay diverging at CPAs and transmission time delay diverging at TSMs [56, 59, 53, 98, 54]. New time delays can be defined for singularities which lack one in the literature. As an example, we will introduce a complex coalescence time delay which diverges at EP-2s in $M=2$ port non-Hermitian scattering systems. Exceptional points occur when $\lambda_{1}-\lambda_{2}=0$, which can be rewritten in terms of matrix invariants $\sqrt{(\text{Tr}S)^{2}-4\text{det}S}=0$. To avoid the square root of a complex number which leads to phase discontinuities, we square both sides:
@@ Appendix E Complex Time Delay - Generalizations
The typical way to define time delay uses a derivative in energy (or equivalently frequency), but the delay operator can be generalized further as a derivative in $p$ where $p$ is any parameter of the scattering matrix [148, 149]. For example, a phase shifter length $\delta L$ for a graph system like in Figs. 1 and 8, or a Hamiltonian perturbation in simulations like in Figs. 5 and 10.
@@
Appendix F Fundamental Singularities Enumerated
Below we provide two tables of many fundamental scattering singularities, Table. 1 listing well studied ones with known applications, while Table. 2 has singularities unexplored in the literature or introduced for the first time in this paper. In both tables, we are considering generic wave systems with either no or minimal inherent symmetries. Otherwise certain singularities, such as symmetric reflection, are not rare events but required phenomena. This is not a complete list since any possible complex scalar formed out of the scattering matrix would have singularities.
Singularity Name Complex Scalar Function That Goes To $00+i$ 0 Phenomena References
Coherent Perfect Absorption (CPA) $\text{det}S$ Absolute absorption of all injected power [55, 39, 56, 57, 58, 48, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71]
Exceptional Point (EP-2) $\lambda_{j}-\lambda_{k}$ Collapse of eigenbasis, square root splitting of eigenvalues [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 91]
Orthogonality Point $\langle R_{j}|R_{k}\rangle$ Orthogonal eigenbasis for a non-unitary $S$-matrix [91]
Reflectionless Scattering Mode (RSM) $S_{xx}$ Zero reflected power at channel $x$ [97, 95, 60, 105, 96, 99, 100]
Transmissionless Scattering Mode (TSM) $S_{xy}$ Zero transmitted power from channel $y$ to channel $x$ [60, 54, 98, 37, 101]
Reciprocal Transmission in Non-Reciprocal Systems $\delta T_{xy}$ Lorentz symmetry in non-reciprocal system [53]
Symmetric Reflection in Generic Systems $\delta R_{xy}$ Reflection symmetry between channels $x$ and $y$ in non-symmetric system [40, 41, 53]
Complex Transmission Time Delay Zero $\tau_{xy}$ Minimal distortion of transmitted pulse [37, 38]
@@ Appendix F Fundamental Singularities Enumerated
Singularity Name Complex Scalar Function That Goes To $00+i$ 0 Phenomena
Other Principal Invariant Zeros $I_{n}=\sum_{1\leq j<k<\cdots<n<M}\lambda_{j}\lambda_{k}\cdots\lambda_{n}$ Uncertain
@@
Complex Wigner-Smith Time Delay Zero $\tau_{WS}$ Uncertain
Complex Reflection Time Delay Zero $\tau_{xx}$ Minimal distortion of reflected pulse
Complex Reflection Time Delay Difference Zero $\tau_{\delta Rxy}$ Uncertain
Complex Transmission Time Delay Difference Zero $\tau_{\delta Txy}$ Uncertain
Any Complex Generalized Delay Zero (any parameter derivative of any function of the $S$-matrix) $\kappa$ Uncertain
Table 2: Table of less explored fundamental singularities of non-unitary scattering matrices. Provided are the name of the singularity, the complex scalar function that goes to zero, and for some a physical consequence. These singularities all exist, but are not known to have application yet.

### v3

@@ Appendix C Singularities in $d=3$ Parameter Space
As stated in Sec. I, while singularities form speckle patterns of points in $d=2$ dimensional parameter space, in $d=3$ they form continuous curves because singularities always have dimension d2d-2. These curves either form closed loops or trail off the edge of the explored parameter space, with no end points visible. Note that this is the same property we see for the d1d-1 dimensional curves in a $d=2$ dimensional parameter space. In $d=3$ space, the d1d-1 dimensional topologically protected structures such as $\text{Re}[\text{det}S]=0$ take the form of surfaces that are either closed or extend beyond the limits of the explored parameter space. In simulations, it can be seen that sometimes d1d-1 structures in $d=2$ space extend out to infinite value of certain parameters (such as loss $\eta$), so we assume the same must be true for d2d-2 structures in higher dimensional spaces. Experimentally, our tunable perturbations are limited within some range, so we cannot verify this assumption.
@@
In Fig. 10, there is a color gradient in frequency but visually every line appears to be a single solid color. Most scattering singularities, CPAs included, do not move much in frequency under perturbations to a graph system in the form of bond length changes. In $\mathcal{D}=2$ and $3$ dimensional cavities there are more pronounced movements in frequency, though still nothing dramatic. Practically, this means it is easier to move multiple different S-matrix zeros to different spots along the real frequency axis than to drag a single zero to different spots along the real frequency axis. For graphs, most frequency movement is on the order of the mean mode spacing or less, which for this graph is about $40$ MHz. An example of this can be seen in Fig. 4(a) of Ref. [erb2025robustwavesplittersbased], which shows a CPA singularity line extending over $60$ MHz, the data coming from the same graph used in Fig. 10.
Appendix D Experimental Ensemble Details
The statistical results presented in this paper were obtained from 31 ensembles of experimental microwave systems, each with a unique combination of the four parameters $\mathcal{D},M,\beta,\eta$. In Table 1, we provide the parameter values for these ensembles, as well as the estimated average number of modes in each realization and the total number of realizations per ensemble. Note that each realization is made up of either $64000$, $80000$, or $100000$ (depending on the capabilities of the network analyzer used) equally spaced frequency points so that we can well characterize the modes and ensure we identify extreme events such as scattering singularities. Each ensemble is therefore over $2.5$ GB of independent scattering matrix data. Details of how the systems are perturbed to create ensembles are discussed in Sec. II.
$\mathcal{D}$ $M$ $\beta$ $\eta$ Values (Estimated Number of Modes) Number of Realizations
1 1 1 2.4 (116), 7.3 (121) 715 for all ensembles
1 1 2 3.1 (119), 7.5 (124) 715 for all ensembles
1 2 1 1.8 (43), 2.1 (65), 2.4 (116), 2.5 (43), 2.8 (65), 4.4 (84), 5.7 (84), 7.0 (84),7.3 (122), 10.1 (84), 11.3 (86), 16.3 (86), 21.4 (86), 30.2 (86) 715 for all ensembles
1 2 2 1.9 (44), 2.9 (44), 3.1 (119), 7.5 (125) 715 for all ensembles
1 3 1 2.5 (44) 715 for all ensembles
1 3 2 3.1 (45), 3.8 (47) 715 for all ensembles
2 2 1 39.0 (68), 40.2 (145), 41.5 (76) 302, 324, 301
2 3 1 40.2 (145) 196
3 2 1 49.0 (1000) 804
3 3 1 49.0 (520), 69.1 (1100), 79.2 (1400), 93.0 (1700), 145 (2100) 804 for all ensembles
Table 1: Table of experimental ensemble parameters. The first three columns correspond to system dimension $\mathcal{D}$, number of channels $M$, and symmetry class $\beta$. Since there are often multiple ensembles that have the same values of $\mathcal{D},M$, and $\beta$, these columns are shared for the sake of compactness. The fourth column has the ensemble value of the cavity absorption $\eta$ (as well as the estimated number of modes in the bandwidth measured). The number of modes is estimated by the bandwidth divided by the mean mode spacing $\Delta$. The final column has the number of realizations in the ensemble.
Appendix E Singularity Statistics - Random Matrix Theory
Figure 11: (a-b) PDFs of $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$ and $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ from 51 RMT ensembles with different values for the three parameters $M,\beta,\eta$. Color gradient of curves corresponds to $\eta$ (not uniformly spaced) of the ensembles. Dashed black lines characterize the power law of the tail behavior. (c-d) PDFs of $\frac{\langle|\text{Re}[\mathcal{S}]|\rangle}{|\text{Re}[\mathcal{S}]|}$ and $\frac{\langle|\mathcal{S}|\rangle}{|\mathcal{S}|}$ for various functions $\mathcal{S}$ from two arbitrary RMT ensembles. Dashed black lines characterize the power law of the tail behavior.
The process used to arrive at statistical results from RMT is as follows. First, we create two sets of $600$ independent Hamiltonians that are of dimension $10^{5}\crossproduct 10^{5}$, generated using GOE (for $\beta=1$) and Gaussian Unitary Ensemble (GUE) (for $\beta=2$) algorithms. Then normalized impedance matrices $Z$ are calculated using the Random Coupling Model [Hemmady2005, Zheng20052Port, Zheng2006, shaibe2025] with the only parameter changed between different ensembles being the dimensionless uniform attenuation $\eta$. The impedance matrices are converted to scattering matrices through the formula:
@@ Appendix E Singularity Statistics - Random Matrix Theory
Note that because the impedance matrices created through RCM are normalized, the characteristic impedance $Z_{0}=1$. The result is a collection of ensembles of $600$ independent $S$-matrices, each ensemble characterized by $M,\beta$ and $\eta$. We treat these scattering matrices generated through RMT the same as our experimental ensembles and conduct the same analysis to calculate singularity mode density, shown in Fig. 4(b).
@@
The RMT data can also be used to calculate PDFs of various complex scalar functions. In Fig. 11, we show the same PDFs as in Fig. 3: (a) $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$, (b) $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$, and the same eight and ten functions in panels (c) and (d), respectively. The simulations utilize two sets of $600$ independent Hamiltonians that are of dimension $10^{5}\crossproduct 10^{5}$, generated using GOE (for $\beta=1$) and Gaussian Unitary Ensemble (GUE) (for $\beta=2$) algorithms. Then the scattering matrices are calculated using the Random Coupling Model [Hemmady2005, Zheng20052Port, Zheng2006, shaibe2025] with the only parameter changed between different ensembles being the dimensionless uniform attenuation $\eta$.
@@
Identical behavior is seen in Fig. 11 as in Fig. 3: 2-2 power law tails are seen panels (a) and (c), and 3-3 power law tails are seen in panels (b) and (d). This shows that even though the data in this paper is measured from microwave cavities, the results we present are true for all chaotic wave scattering systems. The present paper also explains why in the Appendix of Ref. [shaibe2025], a 2-2 power law tail was seen for the probability distributions of complex reflection and transmission time delay differences ($\tau_{\delta R}$ and $\tau_{\delta T}$) from Hermitian ($\eta=0$) systems. As shown in Fig. 5(a), in a Hermitian system, $\delta R_{12}=0$ is d1d-1 dimensional, not a d2d-2 dimensional singularity. The same is true of $\delta T_{12}=0$ if the system is also non-reciprocal. Further the PDFs of complex transmission time delay from a Hermitian, reciprocal system would have 2-2 power laws, which is why the authors of Ref. [shaibe2025] chose to show plots from $\beta=2$ RMT data, since at the time they did not have the topology argument to understand why they did not see 3-3 power laws.
@@
A final comment is that the PDF of $\frac{\langle|C|\rangle}{|C|}$, where $|C|$ is the coalescence of the $S$-matrix eigenvectors, from the reciprocal system shown in panel (c) has the expected 2-2 power law tail out to the largest values, never transitioning to 3-3 as seen in the experimental data in Fig. 3(c). This is because RMT numerics lacks the small amount of residual non-reciprocity we always see in experimental data.
Appendix F Exceptional Points at Infinitesimal Absorption
@@
Figure 12: Phase of $S_{+i}$ and $S_{-i}$ from the same kind of simulation as in Fig. 5 over the same parameter space, but with smaller $\eta=6*10^{-9}$. Black circles surround EPs (phase singularities) and arrows show direction of phase winding.
@@
Appendix G Complex Time Delay - Generalizations
Every singularity can be associated with the divergence of a complex time delay. Examples include Wigner-Smith time delay diverging at CPAs and transmission time delay diverging at TSMs [Lei2020, Erb2024, shaibe2025, Kang2021, Huang2022] , with more examples given in Table. 2. New time delays can be defined for singularities which lack one in the literature. As an example, we will introduce a complex coalescence time delay which diverges at EP-2s in $M=2$ port non-Hermitian scattering systems. Scattering matrix Exceptional points occur when $\lambda_{1}-\lambda_{2}=0$, which can be rewritten in terms of matrix invariants $\sqrt{(\text{Tr}S)^{2}-4\text{det}S}=0$. To avoid the square root of a complex number which leads to phase discontinuities, we square both sides:
@@ Appendix G Complex Time Delay - Generalizations
Singularity Name Complex Scalar Function That Goes To $0+i0$ Phenomena Diverging Complex Time Delay
Coherent Perfect Absorption (CPA) $\text{det}S$ Absolute absorption of all injected power $\tau_{WS}=\frac{-i}{M}\frac{\partial}{\partial E}\text{log}[\text{det}S]$
Exceptional Point (EP-2) $\lambda_{j}-\lambda_{k}$ Collapse of eigenbasis, square root splitting of eigenvalues $\tau_{C}=-i\frac{\partial}{\partial E}\text{log}[(\text{Tr}S)^{2}-4\text{det}S]$
Reflectionless Scattering Mode (RSM) $S_{xx}$ Zero reflected power at channel $x$ $\tau_{xx}=-i\frac{\partial}{\partial E}\text{log}[S_{xx}]$
Transmissionless Scattering Mode (TSM) $S_{xy}$ Zero transmitted power from channel $y$ to channel $x$ $\tau_{xy}=-i\frac{\partial}{\partial E}\text{log}[S_{xy}]$
Reciprocal Transmission in Non-Reciprocal Systems $\delta T_{xy}=S_{xy}-S_{yx}$ Lorentz symmetry in non-reciprocal system $\tau_{\delta T}=-i\frac{\partial}{\partial E}\text{log}[S_{xy}-S_{yx}]$
Symmetric Reflection in Generic Systems $\delta R_{xy}=S_{xx}-S_{yy}$ Reflection symmetry between channels $x$ and $y$ in non-symmetric system $\tau_{\delta R}=-i\frac{\partial}{\partial E}\text{log}[S_{xx}-S_{yy}]$
Table 2: Table of fundamental singularities of non-unitary scattering matrices. Provided are the name of the singularity, the complex scalar function that goes to zero, a physical consequence of the singularity, and the complex time delay that diverges at that singularity.
The typical way to define time delay uses a derivative in energy (or equivalently frequency), but the delay operator can be generalized further as a derivative in $p$ where $p$ is any parameter of the scattering matrix [Ambichl2017, Ambichl2018]. For example, a phase shifter length $\delta L$ for a graph system like in Figs. 1 and 10, or a Hamiltonian perturbation in simulations like in Figs. 5 and 12.
@@
Appendix H Fundamental Singularities Enumerated
Below we provide two tables of many fundamental scattering singularities, Table. 3 listing well studied ones with known applications, while Table. 4 has singularities unexplored in the literature or introduced for the first time in this paper. In both tables, we are considering generic wave systems with either no or minimal inherent symmetries. Otherwise certain singularities, such as symmetric reflection, are not rare events but required phenomena. This is not a complete list since any possible complex scalar function formed out of the scattering matrix would have singularities.
Singularity Name Complex Scalar Function That Goes To $0+i0$ Phenomena References
Coherent Perfect Absorption (CPA) $\text{det}S$ Absolute absorption of all injected power [Chong2010, Fyodorov2017, Lei2020, Imani2020, Frazier2020, Hougne2021, Erb2024, Faul2025, Gupta2012, Pichler2019, Baranov2017, Song2014, Meng2017, IglesiasMartinez2025, Li2022, Vetlugin2021, Vetlugin2022_Resolution, Vetlugin2022_AHOM, erb2025robustwavesplittersbased]
Exceptional Point (EP-2) $\lambda_{j}-\lambda_{k}$ Collapse of eigenbasis, square root splitting of eigenvalues [Berry2004, El-Ganainy18, Alu19, Özdemir19, Ashida20, Ding2022, PhysRevLett.106.213901, PhysRevLett.112.203901, Hodaei2017, Chen2017, PhysRevLett.86.787, PhysRevE.69.056216, PhysRevLett.106.150403, Gao2015, Shi2016, Doppler2016, Stehmann_2004, Suntharalingam2023, erb2025]
Orthogonality Point $\langle R_{j}|R_{k}\rangle$ Orthogonal eigenbasis for a non-unitary $S$-matrix [erb2025]
Reflectionless Scattering Mode (RSM) $S_{xx}$ Zero reflected power at channel $x$ [Jiang2024, Sol2023, Faul2025, Ferise2022, Davy2023, Sweeney2020, Stone2021]
Transmissionless Scattering Mode (TSM) $S_{xy}$ Zero transmitted power from channel $y$ to channel $x$ [Faul2025, Huang2022, Kang2021, Asano2016, Genack2024]
Reciprocal Transmission in Non-Reciprocal Systems $\delta T_{xy}$ Lorentz symmetry in non-reciprocal system [shaibe2025]
Symmetric Reflection in Generic Systems $\delta R_{xy}$ Reflection symmetry between channels $x$ and $y$ in non-symmetric system [Fyodorov2019, Osman2020, shaibe2025]
Complex Transmission Time Delay Zero $\tau_{xy}{\color[rgb]{0,0,0}=-i\frac{\partial}{\partial E}\text{log}[S_{xy}]}$ Minimal distortion of transmitted pulse [Asano2016, giovannelli2025]
@@ Appendix H Fundamental Singularities Enumerated
Singularity Name Complex Scalar Function That Goes To $0+i0$ Phenomena
Other Principal Invariant Zeros of $M\times M$ $S$-matrix $I_{n}=\sum_{1\leq j<k<\cdots<n<M}\lambda_{j}\lambda_{k}\cdots\lambda_{n}$ Uncertain
@@
Complex Wigner-Smith Time Delay Zero $\tau_{WS}{\color[rgb]{0,0,0}=\frac{-i}{M}\frac{\partial}{\partial E}\text{log}[\text{det}S]}$ Uncertain
Complex Reflection Time Delay Zero $\tau_{xx}{\color[rgb]{0,0,0}=-i\frac{\partial}{\partial E}\text{log}[S_{xx}]}$ Minimal distortion of reflected pulse
Complex Reflection Time Delay Difference Zero $\tau_{\delta R}{\color[rgb]{0,0,0}=-i\frac{\partial}{\partial E}\text{log}[S_{xx}-S_{yy}]}$ Uncertain
Complex Transmission Time Delay Difference Zero $\tau_{\delta T}{\color[rgb]{0,0,0}=-i\frac{\partial}{\partial E}\text{log}[S_{xy}-S_{yx}]}$ Uncertain
Any Complex Generalized Delay Zero (any parameter derivative of any function of the $S$-matrix) $\kappa$ (See Eq. 15 for an example) Uncertain
Table 4: Table of less explored fundamental singularities of non-unitary scattering matrices. Provided are the name of the singularity, the complex scalar function that goes to zero, and for some a physical consequence. These singularities all exist, but are not known to have applications yet.